# Discovery potential of a long-lived partner of inelastic dark matter at MATHUSLA in $U(1)_{X_{3}}$ extension of the standard model

### v1

@@ Abstract
We investigate the discovery potential at the MATHUSLA experiment of a long-lived particle (LLP), which is the heavier state of inelastic scalar dark matter (DM) in third generation-philic $U(1)$ ($U(1)_{X_{3}}$) extension of the standard model. Since the heavier state and DM state form the complex scalar charged under the $U(1)_{X_{3}}$, it is natural that the heavier state $P$ is almost degenerate with the DM state and hence long-lived. We find that third generation-philic right-handed $U(1)$, $U(1)_{R_{3}}$, model is the most interesting, because third generation-philic models are less constrained by the current experimental results and right-handed $U(1)$ interactions leave visible final decay products without produing neutrinos. For a benchmark of the model parameters consistent with the current phenomenological constraints, we find that the travel distance of the LLP can be $\mathcal{O}(100)$ m and the LLP production cross section at the 14 TeV LHC can be $\mathcal{O}(10)$ fb. Thus, we conclude that the LLP can be discovered at the MATHUSLA with a sufficiently large number of LLP decay events inside the MATHUSLA detector.
@@ I Introduction
Explanation of the neutrino oscillation phenomena between neutrino flavor species due to non-zero neutrino mass and existence of dark matter (DM) in the Universe require undiscovered particles and interactions of beyond the standard model (BSM) of particle physics. Nevertheless, any evidence of new particles in BSM have not yet been reported in any ongoing experiments. A reason of non-discovery of a new particle might be due to not its large mass but its extremely weak interaction strength, in other words, a relevant coupling constant is very small. Even if coupling constants and hence the resultant production cross sections of new particles are very small, we may expect certain amount of production of new particles if experiments are conducted with huge luminosity. The small coupling constant makes the new particle long-lived, and signals of such long-lived particles (LLPs), if they are electrically neutral, would be discovered through their “displaced vertex” signatures. In fact, displaced vertex signatures have been searched for at, for instance, ATLAS ATLAS:2022pib and FASER FASER:2023tle in Large Hadron Collider (LHC). Some near-future experiments such as FASER2 FASER:2018eoc will start soon, and other several future experiments have been proposed, targeting various range of LLP masses and decay lengths. Among such proposed future experiments for the displaced vertex search, MAssive Timing Hodoscope for Ultra Stable neutraL pArticles (MATHUSLA) experiment is striking, because it will be able to search LLPs with the decay length of the order of $100$Chou:2016lxi . Theoretical studies on search sensitivity for various BSM models have been examined in Refs. BhupalDev:2016nfr ; Evans:2017lvd ; Helo:2018qej ; Deppisch:2018eth ; Jana:2018rdf ; Bauer:2018uxu ; Curtin:2018ees ; Berlin:2018jbm ; Dercks:2018eua ; Deppisch:2019kvs ; No:2019gvl ; Wang:2019xvx ; Jodlowski:2019ycu ; Bolton:2019pcu ; Hirsch:2020klk ; Jana:2020qzn ; Gehrlein:2021hsk ; Sen:2021fha ; Guo:2021vpb ; Bhattacherjee:2021rml ; Du:2021cmt ; Kamada:2021cow ; Bertuzzo:2022ozu ; Liu:2022ugx ; Bandyopadhyay:2022mej ; Mao:2023zzk ; Jodlowski:2023yne ; Fitzpatrick:2023xks ; Curtin:2023skh ; Batz:2023zef ; Deppisch:2023sga ; Bernal:2023coo ; Bishara:2024rtp ; Liu:2024azc ; deVries:2024mla ; Liebersbach:2024kzc .
@@
In the models of Ref. Okada:2019sbb , both DM state and the heavier state relevant for both co-annihilation and inelastic scattering of DM state originate from one complex scalar field. Since they belong to the single complex scalar field, those two states are naturally well degenerate. The decay rate of the heavier state into the lighter one (DM state) would be very small due to its very small phase space volume, so that the heavier state could be a natural candidate for the LLP that we are interested in. To examine the MATHUSLA prospects, inelastic DM models have been proposed Berlin:2018jbm ; Guo:2021vpb ; Bertuzzo:2022ozu  [^1] The mass splitting in Refs. Guo:2021vpb ; Bertuzzo:2022ozu is generated by an explicit gauge symmetry breaking term, and models have not been formulated in gauge invariant manner.. While the longevity of LLP in fermionic inelastic DM model proposed in Ref. Berlin:2018jbm is due to the mass degeneracy as well as its tiny kinetic mixing of dark photon, the LLP in our scalar inelastic DM model is long-lived due to only about $1\%$ mass degeneracy.
This paper is organized as follows: In Sec. II, we describe the extra gauge $U(1)$ interacting inelastic model, which is a variant of the original models in Ref. Okada:2019sbb . We identify a parameter set to reproduce thermal DM abundance and a benchmark point in Sec. III. In Sec. IV, for the benchmark point, we present a search prospect of LLPs by the MATHUSLA experiment. Section V is devoted to our summary.
@@ II The Model
There are several possibilities of anomaly free gauged $U(1)$ extension of the standard model. Among them, the best studied model is the flavor universal $U(1)_{B-L}$ Pati:1973uk ; Davidson:1978pm ; Mohapatra:1980qe ; Mohapatra:1980 . However, because the new neutral gauge boson $Z^{\prime}$ in the $U(1)_{B-L}$ model interacts with the first and the second generation of SM fermions, the experimental constraints on the couplings are very stringent Carena:2004xs ; Amrith:2018yfb ; Das:2019fee . Thus, we may not expect a sizable production cross section for the LLP production from the $Z^{\prime}$ boson decay. Note that anomaly cancellation of $U(1)_{B-L}$ model is realized in each generation of fermions, hence it is generally possible for only a particular generation to be charged under the $U(1)$. The model where only the third generation fermions are charged, $U(1)_{(B-L)_{3}}$, is such a choice Okada:2019sbb ; Babu:2017olk ; Alonso:2017uky ; Bian:2017rpg ; Cox:2017rgn and well studied. As expected, the experimental constraints on the $U(1)_{(B-L)_{3}}$ model are much weaker compared with the universal $U(1)_{B-L}$ model delAmoSanchez:2010bt ; Faroughy:2016osc ; Aaboud:2017sjh ; Chun:2018ibr ; Elahi:2019drj . It is easy to generalize the $U(1)_{B-L}$ model to the $U(1)_{X}$ model by assigning the $U(1)$ charge for a field as a linear combination of its $U(1)_{Y}$ and $U(1)_{B-L}$ charge. Since $U(1)_{Y}$ and $U(1)_{B-L}$ are independently anomaly free, the $U(1)_{X}$ is automatically anomaly free Appelquist:2002mw ; Oda:2015gna ; Das:2016zue . A special case is so-called $U(1)_{R}$ model, where only right-handed (RH) SM fermions and RH neutrinos are charged under the $U(1)$ Jung:2009jz . In this paper, we generalize $U(1)_{(B-L)_{3}}$ to $U(1)_{X_{3}}$, where $X_{3}$ denotes a linear combination of $U(1)_{Y_{3}}$ and $U(1)_{(B-L)_{3}}$. We introduce an $U(1)_{X_{3}}$ charged scalar field $\phi_{1}$ with its charge $+1$ in addition to the minimal particle contents. The total particle contents are listed in Tab. 1. The field $\phi_{2}$ is responsible to break $U(1)_{X_{3}}$ gauge symmetry and $x_{H}$ is a real parameter which parameterize a combination weight of $(B-L)_{3}$ and $Y_{3}$.
@@
where $X^{\mu}$ is the $U(1)_{X_{3}}$ gauge field, $g_{X_{3}}$ is the gauge coupling constant, and $q_{f_{jL/R}}$ is a $U(1)_{X_{3}}$ charge of $f_{jL/R}$ (see Tab. 1).
The scalar potential, which is gauge-invariant and renormalizable, is expressed as Okada:2018xdh ; Chao:2017ilw
@@ II.1 Dark matter mass and interactions
At the $U(1)_{X_{3}}$ and the electroweak (EW) symmetry breaking vacuum, the SM Higgs field and the $U(1)$ Higgs field are expanded around those VEVs, $v$ and $v_{2}$, as (in the unitary gauge)
@@
For a small mixing angle $\alpha$, $h$ is identified with the SM-like Higgs boson. Hereafter, we set $\alpha$ negligibly small. With the $U(1)_{X_{3}}$ and the EW symmetry breaking, the $Z^{\prime}$ boson which is the mass eigenstate after the $X$ becomes massive, $S$ and $P$ acquire their masses, respectively, as
@@
and similarly 3rd generation quarks and leptons also interact with $Z^{\prime}$ boson with corresponding charges. The absence of $Z^{\prime}$-DM-DM coupling indicates that the $Z^{\prime}$-mediating DM scattering off with a nucleon is inelastic and ineffective for a mass splitting larger than the maximal energy transfer in the scatterings Hall:1997ah ; TuckerSmith:2001hy . Elastic scattering through Higgs bosons exchange Jungman:1995df can be neglected since we have set a very small Higgs mixing $\alpha$. We also set the coupling constant $A$ very small, so that $S$ and $P$ are well degenerate in mass. This degeneracy is crucial not only for the scalar $P$ being long-lived, but also for reproducing the observed DM relic density through $S$ and $P$ coannihilation process.
@@ II.2 The decay width of $Z^{\prime}$ boson
where the number of color $N_{c}$ is $3$ for quarks and $1$ for leptons. If $m_{f}\ll m_{Z^{\prime}}$, then Eq. (13) is reduced to
@@
from the vertex (12). The decay branching ratios are shown in Fig. 1. Note that $Br(Z^{\prime}\rightarrow\nu_{\tau}\nu_{\tau})$ vanishes at $x_{H}=-2$, which corresponds to $U(1)$ right-handed, $U(1)_{R}$. This fact plays an important role in the following discussion.
@@ II.3 Decay of $P$
If $S$ and $P$ are strongly degenerate so that the mass difference is much smaller than the mass of $Z^{\prime}$, the two body decay of $P$ is kinematically forbidden. Thus, for the main decay mode, $P\rightarrow SZ^{\prime}{}^{*}\rightarrow Sf\bar{f}$, the total decay width ($\Gamma$), or equivalently the inverse of the lifetime ($\tau$),
@@ III Dark matter and LHC constraints
For evaluation of prospect at MATHUSLA in the next section, in this section, we will find a benchmark point which satisfies the thermal DM abundance and the latest LHC constraints. As metioned above, we concentrate on the $U(1)_{R_{3}}$ model. The LLP $P$ in other general $U(1)_{X_{3}}~{}(x_{H}\neq-2)$ may decay into neutrinos which are invisible for the MATHUSLA detector. Thus, the $U(1)_{R_{3}}$ model is the best choice from the viewpoint of the detection with a large decay length of $\mathcal{O}(100)$ m.
@@ III.1 Thermal relic abundance
Figure 3: The contour (in blue) along which the observed DM relic abundance $\Omega h^{2}\simeq 0.1$ is reproduced for $g_{R_{3}}=0.1$ and $x_{H}=-2$.
@@ III.2 Benchmark points
To find a viable benchmark point, we also take the $Z^{\prime}$ boson search at the LHC with ditau final state CMS:2024pjt ; ATLAS:2019erb into account. We evaluate the production cross section for the process $pp\rightarrow\tau^{+}\tau^{-}$ given by
where the branching ratio is calculated from Eq. (13). As can be seen in Eq. (14), $\mathrm{BR}(Z^{\prime}\rightarrow\tau^{+}\tau^{-})$ depends on the value of $x_{H}$, and we forcus on $x_{H}=-2$ as stated above.
With the decay rates of $X$ estimated in Sec. II.2, as in Ref. Das:2019fee , since the total $X$ boson decay width is very narrow, we use the narrow width approximation to evaluate the $X$ boson production cross section
@@
where $f_{q}$ and $f_{\bar{q}}$ are the parton distribution function (PDF) for a quark and antiquark ($b$ and $\bar{b}$ in our model), $\hat{s}=xys$ is the invariant mass squared of colliding quarks for the center of mass energy $s$. The factor $2$ in Eq. (22) counts two ways of $q$ coming from which proton out of two colliding protons. Since the most severe bound is from the dilepton channel ($\ell=\tau$), we calculate $\sigma(pp\rightarrow X)\mathrm{Br}(X\rightarrow\tau^{+}\tau^{-})$ and compare it with the CMS results CMS:2024pjt  [^2] The ATLAS bound ATLAS:2019erb is somewhat weaker than that from the CMS.. We employ PDFs of CTEQ6L Pumplin:2002vw with a factorization scale $Q=m_{Z^{\prime}}$ for simplicity.
The resultant production cross section is shown in the second column in Tab. 3. The production cross section is consistent with and smaller than the latest most stringent bound from the CMS CMS:2024pjt at the LHC. We summarize parameters and masses of relevant particles in Tab. 2.
@@
$BP_{-2}$ $-2$ $0.1$ $351$ GeV $150$ GeV
Table 2:

### v2

@@ Abstract
We investigate the discovery potential at the MATHUSLA experiment of a long-lived particle (LLP), which is the heavier state of inelastic scalar dark matter (DM) in third-generation-philic $U(1)$ ($U(1)_{X_{3}}$) extension of the Standard Model. Since the heavier state and DM state form the complex scalar charged under the $U(1)_{X_{3}}$, it is natural that the heavier state $P$ is almost degenerate with the DM state and is hence long-lived. We find that third generation-philic right-handed $U(1)$, $U(1)_{R_{3}}$, model is the most interesting, because third-generation-philic models are less constrained by the current experimental results and right-handed $U(1)$ interactions leave visible final decay products without produing neutrinos. For a benchmark of the model parameters consistent with the current phenomenological constraints, we find that the travel distance of the LLP can be $\mathcal{O}(100)$ m and the LLP production cross section at the 14 TeV LHC can be $\mathcal{O}(10)$ fb. Thus, we conclude that the LLP can be discovered at the MATHUSLA with a sufficiently large number of LLP decay events inside the MATHUSLA detector.
@@ I Introduction
Explanation of the neutrino oscillation phenomena between neutrino flavor species due to nonzero neutrino mass and the existence of dark matter (DM) in the Universe require undiscovered particles and interactions of beyond the standard model (BSM) of particle physics. Nevertheless, any evidence of new particles in BSM have not yet been reported in any ongoing experiments. A reason of nondiscovery of a new particle might be due to not their large mass but their extremely weak interaction strength; in other words, the relevant coupling constant is very small. Even if coupling constants and hence the resultant production cross sections of new particles are very small, we may expect a certain amount of production of new particles if experiments are conducted with huge luminosity. The small coupling constant makes the new particle long-lived, and signals of such long-lived particles (LLPs), if they are electrically neutral, would be discovered through their “displaced vertex” signatures. In fact, displaced vertex signatures have been sought at, for instance, ATLAS ATLAS:2022pib and FASER FASER:2023tle in the Large Hadron Collider (LHC). Some near-future experiments such as FASER2 FASER:2018eoc will start soon, and several other future experiments have been proposed, targeting various range of LLP masses and decay lengths. Among such proposed future experiments for the displaced vertex search, MAssive Timing Hodoscope for Ultra Stable neutraL pArticles (MATHUSLA) experiment is striking, because it will be able to search for LLPs with a decay length of the order of $100$Chou:2016lxi . Theoretical studies on the search sensitivity for various BSM models have been examined in Refs. BhupalDev:2016nfr ; Evans:2017lvd ; Helo:2018qej ; Deppisch:2018eth ; Jana:2018rdf ; Bauer:2018uxu ; Curtin:2018ees ; Berlin:2018jbm ; Dercks:2018eua ; Deppisch:2019kvs ; No:2019gvl ; Wang:2019xvx ; Jodlowski:2019ycu ; Bolton:2019pcu ; Hirsch:2020klk ; Jana:2020qzn ; Gehrlein:2021hsk ; Sen:2021fha ; Guo:2021vpb ; Bhattacherjee:2021rml ; Du:2021cmt ; Kamada:2021cow ; Bertuzzo:2022ozu ; Liu:2022ugx ; Bandyopadhyay:2022mej ; Mao:2023zzk ; Jodlowski:2023yne ; Fitzpatrick:2023xks ; Curtin:2023skh ; Batz:2023zef ; Deppisch:2023sga ; Bernal:2023coo ; Bishara:2024rtp ; Liu:2024azc ; deVries:2024mla ; Liebersbach:2024kzc .
@@
In the models of Ref. Okada:2019sbb , both the DM state and the heavier state relevant for both coannihilation and inelastic scattering of the DM state originate from one complex scalar field. Since they belong to a single complex scalar field, those two states are naturally well degenerate. The decay rate of the heavier state into the lighter one (DM state) would be very small due to its very small phase space volume, so that the heavier state could be a natural candidate for the LLP that we are interested in. To examine the MATHUSLA prospects, inelastic DM models have been proposed Berlin:2018jbm ; Guo:2021vpb ; Bertuzzo:2022ozu  [^1] The mass splitting in Refs. Guo:2021vpb ; Bertuzzo:2022ozu is generated by an explicit gauge symmetry breaking term, and models have not been formulated in gauge invariant manner.. While the longevity of LLP in fermionic inelastic DM model proposed in Ref. Berlin:2018jbm is due to the mass degeneracy as well as its tiny kinetic mixing of dark photons, the LLP in our scalar inelastic DM model is long-lived due to only about $1\%$ mass degeneracy.
This paper is organized as follows: In Section II, we describe the extra gauge $U(1)$ interacting inelastic model, which is a variant of the original models in Ref. Okada:2019sbb . We identify a parameter set to reproduce thermal DM abundance and a benchmark point in Section III. In Section IV, for the benchmark point, we present the search prospects of LLPs by the MATHUSLA experiment. Section V is devoted to our summary.
@@ II The Model
There are several possibilities of the anomaly-free gauged $U(1)$ extension of the Standard Model. Among them, the best studied model is the flavor universal $U(1)_{B-L}$ Pati:1973uk ; Davidson:1978pm ; Mohapatra:1980qe ; Mohapatra:1980 . However, because the new neutral gauge boson $Z^{\prime}$ in the $U(1)_{B-L}$ model interacts with the first and the second generation of SM fermions, the experimental constraints on the couplings are very stringent Carena:2004xs ; Amrith:2018yfb ; Das:2019fee . Thus, we may not expect a sizable production cross section for the LLP production from the $Z^{\prime}$ boson decay. Note that anomaly cancellation of the $U(1)_{B-L}$ model is realized in each generation of fermions; hence it is generally possible for only a particular generation to be charged under $U(1)$. A model where only the third-generation fermions are charged, $U(1)_{(B-L)_{3}}$, is such a choice Okada:2019sbb ; Babu:2017olk ; Alonso:2017uky ; Bian:2017rpg ; Cox:2017rgn and has been well studied. As expected, the experimental constraints on the $U(1)_{(B-L)_{3}}$ model are much weaker compared with the universal $U(1)_{B-L}$ model delAmoSanchez:2010bt ; Faroughy:2016osc ; Aaboud:2017sjh ; Chun:2018ibr ; Elahi:2019drj . It is easy to generalize the $U(1)_{B-L}$ model to the $U(1)_{X}$ model by assigning the $U(1)$ charge for a field as a linear combination of its $U(1)_{Y}$ and $U(1)_{B-L}$ charges. Since $U(1)_{Y}$ and $U(1)_{B-L}$ are independently anomaly free, the $U(1)_{X}$ is automatically anomaly free Appelquist:2002mw ; Oda:2015gna ; Das:2016zue . A special case is the so-called $U(1)_{R}$ model, where only right-handed (RH) SM fermions and RH neutrinos are charged under $U(1)$ Jung:2009jz . In this paper, we generalize $U(1)_{(B-L)_{3}}$ to $U(1)_{X_{3}}$, where $X_{3}$ denotes a linear combination of $U(1)_{Y_{3}}$ and $U(1)_{(B-L)_{3}}$. We introduce an $U(1)_{X_{3}}$ charged scalar field $\phi_{1}$ with its charge $+1$ in addition to the minimal particle contents. The total particle contents are listed in Table 1. The field $\phi_{2}$ is responsible for breaking the $U(1)_{X_{3}}$ gauge symmetry and $x_{H}$ is a real parameter that parameterizes a combination weight of $(B-L)_{3}$ and $Y_{3}$. Here, we note that the $U(1)_{(B-L)_{3}}$ and $U(1)_{X_{3}}$ models with a minimal Higgs doublet cannot reproduce realistic fermion masses and flavor mixings due to the gauge symmetry. Since the details of those UV completions are irrelevant for our main discussion on DM and LLP phenomenology, we work on the simplified particle content in Table 1. Nevertheless, we note an example of successful UV completions of $U(1)_{R_{3}}$ in Appendix A222 A few successful UV completions for $U(1)_{(B-L)_{3}}$ can be found in Refs. Babu:2017olk ; Alonso:2017uky ..
@@
where $X^{\mu}$ is the $U(1)_{X_{3}}$ gauge field, $g_{X_{3}}$ is the gauge coupling constant, and $q_{f_{jL/R}}$ is a $U(1)_{X_{3}}$ charge of $f_{jL/R}$ (see Table 1). Here, we do not consider the gauge kinetic mixing because the mixing between $Z$ and $Z^{\prime}$ bosons is experimentally severely constrained; in fact the mixing can also be canceled in a UV model (see Appendix A.). According to the experimental constraints, we have chosen our renormalization condition so as to vanish the kinetic mixing at the scale of $P$ mass.
The scalar potential, which is gauge invariant and renormalizable, is expressed as Okada:2018xdh ; Chao:2017ilw
@@ II.1 Dark matter mass and interactions
At $U(1)_{X_{3}}$ and the electroweak (EW) symmetry breaking vacuum, the SM Higgs field and the $U(1)$ Higgs field are expanded around the VEVs $v$ and $v_{2}$ as (in the unitary gauge)
@@
For a small mixing angle $\alpha$, $h$ is identified with the SM-like Higgs boson. Hereafter, we set $\alpha$ negligibly small. With $U(1)_{X_{3}}$ and the EW symmetry breaking, the $Z^{\prime}$ boson (the mass eigenstate after the $X$ becomes massive), $S$ and $P$ acquire their masses, respectively, as
@@
and similarly third-generation quarks and leptons also interact with the $Z^{\prime}$ boson with corresponding charges. The absence of $Z^{\prime}$-DM-DM coupling indicates that the $Z^{\prime}$-mediating DM scattering with a nucleon is inelastic and ineffective for a mass splitting larger than the maximal energy transfer in the scatterings Hall:1997ah ; TuckerSmith:2001hy . Elastic scattering through Higgs bosons exchange Jungman:1995df can be negligible for a very small Higgs mixing $\alpha$ and the coupling $\lambda_{4}$. We also set the coupling constant $A$ very small, so that $S$ and $P$ are well degenerate in mass. This degeneracy is crucial not only for the scalar $P$ being long-lived, but also for reproducing the observed DM relic density through $S$ and $P$ coannihilation process.
@@ II.2 The decay width of $Z^{\prime}$ boson
where the color number $N_{c}$ is $3$ for quarks and $1$ for leptons. If $m_{f}\ll m_{Z^{\prime}}$, then Eq. (13) is reduced to
@@
from the vertex (12). The decay branching ratios are shown in Fig. 1. Note that $Br(Z^{\prime}\rightarrow\nu_{\tau}\nu_{\tau})$ vanishes at $x_{H}=-2$, which corresponds to right-handed $U(1)$, $U(1)_{R}$. This fact plays an important role in the following discussion.
@@ II.3 Decay of $P$
If $S$ and $P$ are strongly degenerate so that the mass difference is much smaller than the mass of $Z^{\prime}$, the two-body decay of $P$ is kinematically forbidden. Thus, for the main decay mode, $P\rightarrow SZ^{\prime}{}^{*}\rightarrow Sf\bar{f}$, the total decay width ($\Gamma$), or equivalently the inverse of the lifetime ($\tau$),
@@ III Dark matter and LHC constraints
For the evaluation of prospects at MATHUSLA in the next section, in this section, we will find a benchmark point thst satisfies the thermal DM abundance and the latest LHC constraints. As metioned above, we concentrate on the $U(1)_{R_{3}}$ model. The LLP $P$ in other general $U(1)_{X_{3}}~{}(x_{H}\neq-2)$ may decay into neutrinos which are invisible for the MATHUSLA detector. Thus, the $U(1)_{R_{3}}$ model is the best choice from the viewpoint of the detection with a large decay length of $\mathcal{O}(100)$ m.
@@ III.1 Thermal relic abundance
Figure 3: The contour (in blue) along which the observed DM relic abundance $\Omega h^{2}\simeq 0.1$ is reproduced for $g_{R_{3}}=0.1$ and $x_{H}=-2$. Green curves are contours of $\mathrm{Br}(Z^{\prime}\rightarrow SP)$. Here, $g_{R_{3}}=0.1$ in the left pannel and $=0.2$ in the right pannel.
@@ III.2 Benchmark points
To find a viable benchmark point, we take the $Z^{\prime}$ boson search at the LHC into account. With the decay rates of $Z^{\prime}$ estimated in Section II.2, since the total $Z^{\prime}$ boson decay width is very narrow, as in Ref. Das:2019fee , we use the narrow width approximation to evaluate the $Z^{\prime}$ boson production cross section
@@
where $f_{q}$ and $f_{\bar{q}}$ are the parton distribution function (PDF) for a quark and antiquark ($b$ and $\bar{b}$ in our model), $\hat{s}=xys$ is the invariant mass squared of colliding quarks for the center of mass energy $s$. The factor $2$ in Eq. (21) accounts for two ways of $q$ coming from which proton out of two colliding protons. We employ PDFs of CTEQ6L Pumplin:2002vw with a factorization scale $Q=m_{Z^{\prime}}$ for simplicity.
Since the severest bound on the $Z^{\prime}$ boson is from the dilepton channel ($\ell=\tau$CMS:2024pjt ; ATLAS:2019erb 333 We note that the present experimental bound from the $Z^{\prime}\rightarrow Zh$ mode reported by the ATLAS ATLAS:2022enb and CMS CMS:2024phk Collaborations for $m_{Z^{\prime}}=\mathcal{O}(100)$ GeV are about one order of magnitude weaker than that from the $\tau^{+}\tau^{-}$ mode., we evaluate the production cross section for the process $pp\rightarrow\tau^{+}\tau^{-}$ given by
where the branching ratio is calculated from Eq. (13), and compare it with the CMS results CMS:2024pjt  [^4] The ATLAS bound ATLAS:2019erb is somewhat weaker than that from the CMS.. The resultant production cross section is shown in the second column in Table 3. The production cross section is consistent with and smaller than the latest most stringent bound from the CMS CMS:2024pjt at the LHC. We summarize parameters and masses of relevant particles of benchmark points with different values of the $U(1)_{R_{3}}$ gauge coupling constant in Table 2.
@@
$BP_{1}$ $-2$ $0.1$ $351$ GeV $150$ GeV
$BP_{2}$ $-2$ $0.2$ $850$ GeV $351$ GeV
Table 2: The values of parameters and the masses of $Z^{\prime}$ and $S$ in benchmark points.